For IBM, find the following items (Hint: Try using Yahoo Finance & FINRA) which are required for you to calculate the Weighted Average Cost of Capital (WACC):

  • Shares outstanding = 925.794

  • Stock price =             $154.76
  • Equity =                     $16792000
  • Debt =                        $39079250

  • Preferred stock         $N/A

  • Equity beta                1.06%

  • Market risk premium = 5%

  • US treasury yield (10-year) = 2.37%

  • Cost of equity (CAPM)       7.67...%

  • Pre-tax cost of debt = 2.25%      

  • Tax rate                                  …………………………..%

  • Credit rating                          …………………………

  • Cost of preferred stock        …………………………%
  • WACC =                               ………………………………..%
Answer #1

Based on given information we calculate WACC as under other answers are highlighted in bold.

WACC = (E / V) × Re + (D / V) × Rd × (1 − Tc)
(E/V)*Re (D/V)*Rd (1-Tc) WACC
0.023052042 + 0.015738 X 0.79 = 3.55%
E $      16,792,000
D $      39,079,250
V = E+D= $          55,871,250
Re= 7.67%
Rd 2.25%
T= 21% (US Corporate tax rate)
Re = Risk-Free Rate of Return + Beta * (Market Rate of Return – Risk-free Rate of Return)
Equity Beta= 1.06%

Tax rate = 21% -Assumed for US Corpoations (IBM)

Credit Rating= AAA/Sovereign rating-IBM (Not relevant as cost of debt is provided)